Brushy Mountain Mining Company's coal reserves are being depleted, so its sales are falling. Also, environmental costs increase each year, so its costs are rising. As a result, the company's earnings and dividends are declining at the constant rate of 9% per year. If D0 = $2 and rs = 10%, what is the estimated value of Brushy Mountain's stock? Round your answer to the nearest cent.
Answer #1

Current price=D1/(Required return-Growth rate)

=(2*(1-0.09))/(0.1-(0.09)

=1.82/0.19

=$9.58(Approx).